The spectroscopic data given in the table was generated for two different concentrations of a solution. Concentration Absorbance 0.0149 0.1301 0.1279 0.6013 What is the slope of the line formed by these points? What is the ‑intercept? What is the concentration of a solution with an absorbance of ? concentration:
Answer #1

we know, b= A-(mc) I cone Absorbane From the tabler 100149 1 0,1279 10.13011 0.6013 b=0.1979 - 10.01498m)-20 b=0.6013.-60.130